OpenStudy (anonymous):

The midpoint of CD is E(–1, 0). One endpoint is C (5, 2). What are the coordinates of the other endpoint?

OpenStudy (anonymous):

from 5 to -1 is 6 units left go another 6 units left, get to -7

OpenStudy (anonymous):

similarly, from 2 to 0 is 2 units down another 2 units down, you are at -2 so other endpoint is \((-7,-2)\)

OpenStudy (anonymous):

okay thxx that was pretty simple lol

OpenStudy (anonymous):

yeah simpler than solving \[\frac{x+5}{2}=-1\] and \[\frac{y+2}{2}=0\]but you can do it that way too if you like

OpenStudy (anonymous):

yea it is haha , thx

OpenStudy (anonymous):

yw